According to a 2009 Reader's Digest article, people throw away approximately 12% of what they buy at the grocery store. Assume this is the true proportion and you plan to randomly survey 89 grocery shoppers to investigate their behavior. What is the probability that the sample proportion exceeds 0.2?

Answer =______ (Enter your answer as a number accurate to 4 decimal places.)
Answer #1

Solution:

The probability that the sample proportion exceeds 0.2 is 0.0099

Explanation:

P(\bar{p}>0.2)=P\begin{pmatrix} \frac{\bar{p}-\hat{p}}{\sqrt{\frac{\hat{p}(1-\hat{p})}{n}}} >\frac{0.2-0.12}{\sqrt{\frac{0.12(1-0.12)}{89}}} \end{pmatrix}

P(\bar{p}>0.2)=P\begin{pmatrix} Z >2.33 \end{pmatrix}=0.0099

P(\bar{p}>0.2)=0.0099

The probability that the sample proportion exceeds 0.2 is 0.0099

Please rating the answer. Thank you...!!